# Cross Domain Early Crop Mapping using CropGAN and CNN Classifier

### v1

@@ Abstract
Driven by abundant satellite imagery, machine learning-based approaches have recently been promoted to generate high-resolution crop cultivation maps to support many agricultural applications. One of the major challenges faced by these approaches is the limited availability of ground truth labels. In the absence of ground truth, existing work usually adopts the ”direct transfer strategy” that trains a classifier using historical labels collected from other regions and then applies the trained model to the target region. Unfortunately, the spectral features of crops exhibit inter-region and inter-annual variability due to changes in soil composition, climate conditions, and crop progress, the resultant models perform poorly on new and unseen regions or years. This paper presents the Crop Generative Adversarial Network (CropGAN) to address the above cross-domain issue. Our approach does not need labels from the target domain. Instead, it learns a mapping function to transform the spectral features of the target domain to the source domain (with labels) while preserving their local structure. The classifier trained by the source domain data can be directly applied to the transformed data to produce high-accuracy early crop maps of the target domain. Comprehensive experiments across various regions and years demonstrate the benefits and effectiveness of the proposed approach. Compared with the widely adopted direct transfer strategy, the F1 score after applying the proposed CropGAN is improved by $\bf{13.13\%-50.98\%}$.
@@
Early Crop Mapping, Cross Domain, Domain Adaptation, CropGAN, Cropland Data Layer.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Early crop mapping, i.e. determining the cultivation regions of crops before their harvest season, is the fundamental building block for agricultural planning, resource allocation, crop insurance, risk management[1, 2] and many other agricultural applications. Since the spectral features of vegetation are determined by their structure, leaf biochemistry and phonological stages [3], time-series analysis on multi-spectral images captured by satellites is the dominant approach for land cover classifications. With the rapid growth in data volume and increasing accessibility of satellite imagery, deep learning (DL) based remote sensing has been promoted in recent years to produce high-resolution, high-accuracy crop cultivation maps [4]. This class of approaches relies on a large amount of ground truth data, also known as labels, to train and validate the classification model. The ground truth can be obtained from surveys as first-hand labels[5, 6], or use public datasets, such as the United States Department of Agriculture (USDA)’s Cropland Data Layer (CDL), as weak labels for model training [7]. Different DL architectures, such as convolutional neural networks (CNNs), deep autoencoders and recurrent neural networks with long short-term memory (LSTM), are explored for crop mapping tasks [8, 9, 10]. The results suggest DL approaches outperform conventional support vector machine (SVM) and tree-based models in providing semantic information on the input images.
@@
Unfortunately, obtaining appropriate ground truth for an arbitrary region is challenging. Public datasets such as CDL are only available for a few countries and are usually released after harvest season. On the other hand, the training datasets collected by surveys can provide reliable and timely references for model training. Still, the process can be costly, labour-intensive, and sometimes unfeasible, especially in underdeveloped countries. In the absence of ground truth, existing work usually adopts the ”direct transfer strategy” that trains a classifier first using available labelled data for other regions and then applies the trained model to the target region [11, 12]. However, spectral features of crops have both inter-region variability and inter-annual variability due to changes in soil composition, climate conditions, and crop progress [13]. These variability collectively contribute to the distribution shift between the training data (source domain) and the test data (target domain). Consequently, the direct transfer strategy often leads to poor performance on new and unseen regions and years as it compromises an implicit assumption of the machine learning-based crop mapping approaches: the labelled training data and the data from the target region are independent and identically distributed (i.i.d), or at least come from similar distributions.
@@
To address this issue, various methods have been developed aiming to enhance the model’s ability on unseen domains. One effective approach involves training the model by incorporating multi-year data and the respective phonological metrics as the major inputs [14]. This method helps the model to capture the temporal variations in spectral patterns for target crops caused by changing environmental and climate conditions. However, their approach only enhances the model’s ability to generalise over different years of the same region, while the cross-region issue remains unsolved. Transfer learning [11, 12, 15, 16, 17] is a potent method that harnesses knowledge acquired from a source domain to enhance performance in a target domain. This technique involves utilizing pre-trained models and fine-tuning them on the target domain using a small amount of labelled target domain data. By doing so, the model can swiftly adapt to new data distributions and achieve improved performance. This approach works well both on cross-region and cross-year scenarios. Yet, although the amount of labelled data required by fine-tuning a pre-trained model is much smaller than training a model from scratch, collecting labels for the target domain is still labour-intensive and time-consuming.
This paper presents the Crop Generative Adversarial Network (CropGAN) to address domain shift issues that arise in early crop mapping applications, both across regions and years. We assume that only one source domain has labelled data for model training. The objective is to map a specific type of crop (e.g., corn) in a target domain for which no labelled data is available. Toward this requirement, the proposed approach considers applying an unsupervised domain adaptation technique to learn a mapping function to transform the spectral features of the target domain to the source domain while preserving their local structure. For example, the spectral features of corn cultivated in the target domain might exhibit distinct patterns from those in the source domain. After applying CropGAN, the difference between corn in the two domains should be minimized while the corn and other types of land covers in the target domain are still distinguishable. This transformation facilitates knowledge transformation from the labelled source domain to the unlabeled target domain, allowing the target domain to directly use the models trained on the source domain without degrading the performance.
From the highest level, the proposed approach consists of three key components: the pre-processor, the domain mapper, and the crop mapper. The pre-processor module employs linear interpolation to fill gaps due to cloud coverage in the Multi-Spectral Images (MSI), ensuring a complete time series. The domain mapper transforms the time-series MSI remote sensing data from the target domain to the source domain. This is achieved by utilizing an adversarial architecture [18] consisting of two generator networks and two discriminator networks. The generator learns to transform data points from the target domain to the source domain, while the discriminator distinguishes between the transformed data points and the original data points from the source domain. To train such a network, paired data that associate samples of the same category from distinct domains is required. However, the lack of labelled data within the targeted domain impedes the establishment of direct associations between datasets, presenting a barrier to forming explicit correlations between disparate domains. To overcome this obstacle, the domain mapper incorporates a cycle consistency loss [18] to ensure the transformed data points can be accurately reconstructed back into their original domain, eliminating the need for paired data. Finally, the crop mapper, which is a CNN-based model trained on the source domain, can be directly applied to the transformed data to accurately determine the cultivation locations of the specified target crop on the target domain.
@@
  • Propose CropGAN to address the domain shift issue due to the inter-region and inter-year variations in remote sensing based crop mapping. By employing unsupervised domain adaptation techniques with the integration of time-series satellite imagery, the proposed approach effectively addresses the challenges posed by limited or unavailable labelled data in target domains. This approach enables accurate crop mapping in such regions.
  • Design and implement a fully functional system based on the CropGAN approach to accurately locate the specific crop of the interest in regions without labelled data.
  • Conduct cross-region and cross-year experiments using datasets of regions from three countries, the USA, China, and Canada, to evaluate the proposed approach. The results demonstrate superior performance compared to the baseline method, confirming the effectiveness and accuracy of the CropGAN approach for crop mapping.
@@ II Related Works
Remote sensing-based crop-type mapping has undergone significant advancements with the adoption of machine learning methods. Traditionally, techniques such as SVM and random forest (RF) have been widely utilized for crop classification using remote sensing data [19, 20, 21]. However, the emergence of deep learning methods has brought about a revolution in this field by leveraging their ability to automatically extract meaningful representations from data.
@@
Deep learning models, such as LSTM networks and CNNs, have exhibited impressive performance in crop type mapping using remote sensing imagery [8, 10, 22]. They excel at capturing temporal and spatial characteristics, enabling accurate classification of various crop types. By training on labelled crop datasets, these models can learn complex relationships between spectral, spatial, and temporal features, leading to improved accuracy in distinguishing crop types. However, deep learning models encounter an adaptation problem when it comes to effectively capturing and comprehending the intricate relationships and variations in crop patterns across different regions and time periods.
To address this challenge, training the model with multi-year crop data proves to be a valuable adaptation method. Recent studies, including [14], emphasize the importance of incorporating multi-year crop data from a specific region as it enhances the model’s understanding of temporal patterns, variations, and trends within that geographic context. This approach enhances the accuracy and reliability of crop classification by considering interannual climate variations and capturing the unique characteristics of the region. Leveraging region-specific multi-year data enables the model to become more robust to phenology shifts, ensuring consistent performance across different years. Training on multi-year crop data from the same region serves as an effective adaptation technique, resulting in more precise and dependable crop mapping outcomes within a particular geographic area. A key drawback of the multi-year crop data training method is its high reliance on labelled data. To effectively train deep learning models, a large number of accurately labelled samples are needed. However, collecting such a vast amount of labelled data from diverse regions and years is a challenging task. The process is time-consuming, costly, and often impractical due to the effort and resources required.
To address this challenge, transfer learning techniques have been introduced, allowing the knowledge learned from a source domain, where labelled samples are abundant, to be transferred and applied to a target domain with limited labelled samples [11, 15].
@@
Direct transfer [11, 12] is a commonly used transfer learning approach in remote sensing-based crop type mapping. It involves training models on regions with abundant labelled samples and directly applying them to other regions. This approach assumes that the knowledge learned from the source domain is applicable to the target domain. While direct transfer can be a simple and effective method, it may encounter difficulties due to the domain shift problem, where the characteristics of the target region, such as soil types, climate conditions, or crop varieties, differ significantly from those of the source region.
To overcome the limitations of direct transfer and address domain shift, fine-tuning has emerged as a valuable transfer learning technique. Fine-tuning involves taking a pre-trained model from the source domain and further training it using a smaller set of labelled samples from the target domain. By leveraging the knowledge already acquired from the source domain, fine-tuning enables the model to adapt and specialize for the crop classification task in the target domain [23, 24]. This adjustment of the pre-trained model’s parameters allows it to capture the specific features and characteristics of the target domain, enhancing its performance in crop classification.
Fine-tuning has demonstrated promising results in improving the accuracy of crop classification models. For instance, researchers have employed pre-trained deep learning models, such as VGG16 or ResNet, which are initially trained on large-scale datasets like ImageNet, and then fine-tuned them using crop samples from the target domain [16, 17]. By transferring the knowledge acquired from a large and diverse dataset, the fine-tuned models exhibit enhanced capabilities in classifying crops accurately, even with limited labelled samples from the target domain.
Despite the effectiveness of fine-tuning, it still relies on labelled samples from the target domain, which can be a labour-intensive and time-consuming process, especially in developing or undeveloped regions. Compared with the aforementioned supervised methods, unsupervised domain adaptation (UDA) methods can transfer knowledge learned from a source domain with a large number of labelled training samples to a target domain with only unlabeled data [25]. The core idea behind UDA is to align the feature distributions between the source and target domains, ensuring that the model learns transferable representations. By achieving distribution alignment, UDA facilitates effective knowledge transfer and improves the performance of crop classification models in the absence of labelled samples from the target domain. By aligning the feature distributions, these methods enable the model to learn domain-invariant representations, enhancing its ability to generalize from the source domain to the target domain.
Various techniques have been developed for unsupervised domain adaptation in remote sensing-based crop type mapping based on UDA. [26] introduced the deep adaptation crop classification network (DACCN) based on UDA principles for crop type mapping. Additionally, the concept of multi-source UDA (MUDA) has been explored, leveraging knowledge from multiple source domains to achieve superior results compared to single-source UDA (SUDA). [27] proposed the multi-source UDA crop classification model (MUCCM) for unsupervised crop mapping. These UDA approaches alleviate the reliance on labelled data from the target domain, making crop mapping more feasible in various settings.
In this paper, an alternative approach is proposed to address the domain shift problem in early crop mapping without relying on labelled data from the target domain. Instead of focusing on learning transferable representations like UDA, our method employs CropGAN to transform the original unlabelled data from the target domain into the source domain. By leveraging the power of CropGAN, the objective is to bridge the domain gap and enable the model to effectively utilize the labelled data from the source domain for early crop mapping in the target domain. This approach provides a practical solution for overcoming the scarcity of labelled data in the target domain, making crop mapping more feasible and accurate in diverse settings.
@@ III-A Study area
Figure 1: The Study areas for the cross-region problem. (a) Cavalier County, North Dakota, USA. (b) Sussex County, Delaware, USA. (c) Region (from ${}^{\circ}$ to ${}^{\circ}$ latitude and from ${}^{\circ}$ to ${}^{\circ}$ longitude), Jilin Province, China. (d) Region (from ${}^{\circ}$ to ${}^{\circ}$ latitude and from ${}^{\circ}$ to ${}^{\circ}$ longitude), Quebec, Canada.
Figure 2: The Study area for the cross-year problem. Jackson County, Minnesota, USA.
This paper chose corn as the example crop to demonstrate the cross-domain capability of the proposed approach. That is, the objective is to map corn cultivation locations in a target domain using models trained by labelled data from a source domain. The target and source domains could differ in their geographic locations (i.e., cross-region) or be in the same region in different years (i.e., cross-year). Note that, the proposed approach can be applied to other types of crops. Toward this end, the study areas include the following regions: Jackon County, Minnesota, Cavalier County, North Dakota, and Sussex County, Delaware of the USA, and two sub-areas from Jilin Province, China, and Quebec, Canada, respectively. For a visual depiction of the geographical locations of these counties and areas, please refer to Figure 1 and 2.
TABLE I: The environmental conditions of the source domain and the target domains. ID 1 represents Jackson County. ID 2 represents the cavalier county. ID 3 represents the study area in China. ID 4 represents Sussex County. ID 5 represents the study area in Canada. ”T” represents the yearly average temperature. ”P” stands for the average hourly precipitation. ”E” signifies the average hourly evaporation. ”R” indicates the surface net solar radiation.
ID Year T(K) P(mm/h) E(mm/h) R(kJ/m^2)
1 2019 289.02 2.60 -1.01 4688.12
1 2020 289.93 1.48 -1.11 5118.46
1 2021 290.76 1.06 -1.06 5245.53
1 2022 289.66 1.54 -1.11 4967.67
2 2019 286.24 1.04 -0.92 4996.68
3 2019 290.03 1.99 -1.99 12542.54
4 2019 294.37 1.62 -1.29 5962.68
5 2019 288.86 1.27 -1.03 5431.65
Figure 3: The crop calendar for corn. The red lines represent the average dates of the planting periods.
@@
Figure 4: The average time-series NDVI of corn for the source domain and target domains in the USA. (a) Jackson County 2019 vs. Jackson County 2020. (b) Jackson County 2019 vs. Jackson County 2021. (c) Jackson County 2019 vs. Jackson County 2022.
These counties and areas exhibit distinct environmental conditions, including variations in temperature, precipitation, evaluation, and solar radiation. The environmental condition metrics for the selected areas are shown in Table I. Moreover, the crop calendar for corn differs across these areas, as shown in Figure 3. The above differences collectively affect the development of corn in these regions. For example, figure 4a illustrates a notable disparity between the NDVI values observed in the early growth phase of corn in Jackson County between 2019 and 2020. Notably, the NDVI value curve exhibited an earlier increase in 2020 compared to the pattern observed in 2019. Consequently, a model trained on the data from Jackson County in 2019 can not work well in the case of 2020, as the distributions of their spectral features over time are not i.i.d.. Figure 5 complements the observation by comparing the NDVI values of corn between the study areas in the USA and those in China and Canada. In China, the maximum NDVI value is higher than that observed in the USA. Additionally, in Canada, there is a notable trend of increased NDVI values occurring approximately 30 days earlier compared to the pattern observed in the USA.
Figure 5: The average time-series NDVI of corn for the source domain and target domains between USA, China, and Canada. (a) Cavalier County 2019 vs. Study area in China 2019. (b) Sussex County 2019 vs. Study area in Canada 2019.

### v2

# Cross Domain Early Crop Mapping using CropSTGAN

@@ Abstract
Driven by abundant satellite imagery, machine learning-based approaches have recently been promoted to generate high-resolution crop cultivation maps to support many agricultural applications. One of the major challenges faced by these approaches is the limited availability of ground truth labels. In the absence of ground truth, existing work usually adopts the ”direct transfer strategy” that trains a classifier using historical labels collected from other regions and then applies the trained model to the target region. Unfortunately, the spectral features of crops exhibit inter-region and inter-annual variability due to changes in soil composition, climate conditions, and crop progress, the resultant models perform poorly on new and unseen regions or years. Despite recent efforts, such as the application of the deep adaptation neural network (DANN) model structure in the deep adaptation crop classification network (DACCN), to tackle the above cross-domain challenges, their effectiveness diminishes significantly when there is a large dissimilarity between the source and target regions. This paper introduces the Crop Mapping Spectral-temporal Generative Adversarial Neural Network (CropSTGAN), a novel solution for cross-domain challenges, that doesn’t require target domain labels. CropSTGAN learns to transform the target domain’s spectral features to those of the source domain, effectively bridging large dissimilarities. Additionally, it employs an identity loss to maintain the intrinsic local structure of the data. Comprehensive experiments across various regions and years demonstrate the benefits and effectiveness of the proposed approach. In experiments, CropSTGAN is benchmarked against various state-of-the-art (SOTA) methods. Notably, CropSTGAN significantly outperforms these methods in scenarios with large data distribution dissimilarities between the target and source domains.
@@
Early Crop Mapping, Multispectral Image Data, Cross Domain, Domain Adaptation, CropSTGAN, Cropland Data Layer.
[…]
=-21pt
@@ I Introduction
Early crop mapping, i.e. determining the cultivation regions of crops before their harvest season, is the fundamental building block for agricultural planning, resource allocation, crop insurance, risk management[1, 2] and many other agricultural applications. Since the spectral features of vegetation are determined by their structure, leaf biochemistry and phonological stages [3], time-series analysis on multi-spectral images captured by satellites is the dominant approach for land cover classifications. With the rapid growth in data volume and increasing accessibility of satellite imagery, deep learning (DL) based remote sensing has been promoted in recent years to produce high-resolution, high-accuracy crop cultivation maps [4]. This class of approaches relies on a large amount of ground truth data, also known as labels, to train and validate the classification model. The ground truth can be obtained from surveys as first-hand labels[5, 6], or use public datasets, such as the United States Department of Agriculture (USDA)’s Cropland Data Layer (CDL), as weak labels for model training [7]. Different DL architectures, such as convolutional neural networks (CNNs), Temporal Convolutional Neural Network (TempCNN) [8], deep autoencoders and recurrent neural networks with long short-term memory (LSTM), are explored for crop mapping tasks [9, 10, 11]. The results suggest DL approaches outperform conventional support vector machine (SVM) and tree-based models in providing semantic information on the input images.
@@
Unfortunately, obtaining appropriate ground truth for an arbitrary region is challenging. Public crop-type ground truth datasets, such as CDL, provide reliable and timely references for model training. However, these datasets are only available for a few countries and are usually released after harvest season. The collecting process can be costly, labour-intensive, and sometimes unfeasible, especially in underdeveloped countries. In the absence of ground truth, existing work usually adopts the ”direct transfer strategy” that trains a classifier first using available labelled data for other regions and then applies the trained model to the target region [12, 13]. However, spectral features of crops have both inter-region variability and inter-annual variability due to changes in soil composition, climate conditions, and crop progress [14]. These variability collectively contribute to the distribution shift between the training data (source domain) and the test data (target domain), called the cross-domain issue. Consequently, the direct transfer strategy often leads to poor performance on new and unseen regions and years as it compromises an implicit assumption of the machine learning-based crop mapping approaches: the labelled training data from the source domain and the data from the target region are independent and identically distributed (i.i.d), or at least come from similar distributions.
@@
To address this cross-domain issue, including cross-region and cross-year issues, various methods have been developed aiming to enhance the model’s ability on unseen domains. One effective approach involves training the model by incorporating multi-year data and the respective phonological metrics as the major inputs [15]. This method helps the model to capture the temporal variations in spectral patterns for target crops caused by changing environmental and climate conditions. Their methodology augments the model’s proficiency in generalizing across diverse years within a singular region for which multi-annual ground truth data exists. However, the predicament of cross-regional issue remains unaddressed.
To mitigate the challenges associated with both inter-regional and inter-annual cross-domain issues, methodologies such as Domain Adversarial Neural Networks (DANN) [16] and their derivatives, including Self-Training with Domain Adversarial Network (STDAN) [17], Phenology Alignment Network (PAN) [18], and Deep Adaptation Crop Classification Network (DACCN) [19], have been deployed. These approaches endeavour to delineate invariant features from data across both target and source domains, leveraging these features for enhanced accuracy in crop classification mappings. However, these methodologies hinge on the presupposition that the data distributions between the target and source domains exhibit relatively minor disparities. Furthermore, all these methodologies operate under the assumption that the crop types are identical in both the source and target domains. In the STDAN study [17], two case studies comprising four areas from Gangwon Province and Gyeongsang Province, Korea were selected. These areas, while not geographically distant, exhibit slight environmental differences. Similarly, the PAN study [18] chose three areas from Sichuan Province, Hubei Province, and Anhui Province in China, all positioned at similar longitudes, also leading to minor environmental variations (Wang, 2021). The DACCN study [19], addressing cross-country issues for the first time in crop mapping experiments, selected sub-areas from Heilongjiang Province and Jilin Province, China as target domains, with four sub-areas from different states in the USA as source domains. This cross-country setup introduced significantly larger environmental variations compared to the previous studies. Additionally, the study included cross-year experiments in the USA. Notably, DACCN’s performance metrics were significantly lower for cross-country issues than for cross-year experiments, indicating a reduction in efficiency due to substantial discrepancies in data distribution across countries. In other words, although these approaches adeptly tackle the issue of missing labels and diminish the detrimental impacts of domain shifts, their efficacy is constrained by relevantly substantial discrepancies in data distribution across the domains.
To address this domain shift challenge encountered in early crop mapping endeavors under relevantly substantial discrepancies in data distribution across the domains, this paper introduces the Crop Mapping Spectral-temporal Generative Adversarial Neural Network (CropSTGAN) framework. The system’s primary goal is to identify a specified crop variety, such as corn, within a target area lacking labelled data. To fulfil this aim, the CropSTGAN framework employs an unsupervised domain adaptation (UDA) strategy, a technique for adapting a model from a source domain with labelled data to an unlabelled target domain. The innovative method learns a function that transforms the spectral characteristics features of the target area to the source domain while retaining their local structure. For instance, the spectral features of corn in the target domain may diverge from those in the source domain. The CropSTGAN framework first transforms the target domain’s spectral features to resemble those of the source domain, thereby minimizing the discrepancies in feature patterns between the transformed target data and the original source data, all the while maintaining the distinguishability of corn and other land cover types within the target region. This process enables the straightforward application of a crop mapping model, trained within the source domain, to the target domain without diminishing its precision.
From the highest level, the proposed CropSTGAN system consists of three key components: the pre-processor, the CropSTGAN domain mapper, and the TempCNN crop mapper. The pre-processor module employs linear interpolation to fill gaps due to cloud coverage in the Multi-Spectral Images (MSI), ensuring a complete time series. The CropSTGAN domain mapper, a modification of the CycleGAN architecture [20], is designed with a specific structure to capture the temporal features from the time-series MSI data and transform time-series MSI data points from the target domain to the source domain. The CropSTGAN consists of two generator networks and two discriminator networks. The generators learn to transform data points from one domain to another domain, while the discriminators distinguish between the transformed data points provided by the generators and the original data points. Finally, the TempCNN crop mapper, a CNN-based model shared structure with the CropSTGAN discriminators, can be directly applied to the transformed target data to accurately determine the cultivation locations of the specified target crop on the target domain. To evaluate the distinct architecture of the CropSTGAN domain mapper, comparative experiments were undertaken utilizing a simpler, analogous structure named CropTGAN, alongside benchmarking against several state-of-the-art (SOTA) methodologies, including TempCNN and STDAN. The empirical findings corroborated that the CropSTGAN architecture enhances crop mapping accuracy under the relevantly substantial discrepancies in data distribution across the domains.
@@
  • Propose the CropSTGAN framework, consisting of a pre-processor, a CropSTGAN domain mapper, and a TempCNN crop mapper, to address the cross-domain issue due to the inter-region and inter-year variations in remote sensing-based early crop mapping under substantial discrepancies in data distribution across the domains.
  • Design the CropSTGAN domain mapper to capture the temporal and spectral features from the time-series MSI data and learn to transform the target domain data into the source domain.
  • Conduct cross-region and cross-year experiments in the study areas from the USA and China to evaluate the CropSTGAN framework. The results demonstrate superior performance compared to CropTGAN and several SOTA methods, such as TempCNN and STDAN, confirming the effectiveness and accuracy of the CropSTGAN framework for cross-domain early crop mapping.
@@ II Related Works
Remote sensing-based crop-type mapping has undergone significant advancements with the adoption of machine learning methods. Traditionally, techniques such as SVM and random forest (RF) have been widely utilized for crop classification using remote sensing data [21, 22, 23]. However, the advent of deep learning has brought about a revolution in this field by leveraging their ability to automatically extract meaningful representations from data.
@@
Deep learning models, such as LSTM networks and CNNs, have exhibited impressive performance in crop type mapping using remote sensing imagery [9, 11, 24]. They excel at capturing temporal and spatial characteristics, enabling accurate classification of various crop types. By training on labelled crop datasets, these models can learn complex relationships between spectral, spatial, and temporal features, leading to improved accuracy in distinguishing crop types.
@@
To locate the target crop in the target domain without ground truth data, direct transfer techniques have been introduced, allowing the knowledge learned from a source domain, where labelled samples are abundant, to be directly applied to the target domain [12, 25]. It involves training models on regions with abundant labelled samples and directly applying them to other regions or years. This approach assumes that the knowledge learned from the source domain is applicable to the target domain. While direct transfer can be a simple and effective method, it may encounter a cross-domain problem when it comes to effectively capturing and comprehending the intricate relationships and variations in crop patterns across different regions and years.
To address the cross-year issue, training the model with multi-year crop data proves to be a valuable adaptation method. Recent studies, including [15], emphasize the importance of incorporating multi-year crop data from a specific region as it enhances the model’s understanding of temporal patterns, variations, and trends within that geographic context. This approach enhances the accuracy and reliability of crop classification by considering inter-annual climate variations and capturing the unique characteristics of the region. Leveraging region-specific multi-year data enables the model to become more robust to phenology shifts, ensuring consistent performance across different years. Training on multi-year crop data from the same region serves as an effective adaptation technique, resulting in more precise and dependable crop mapping outcomes within a particular geographic area. A key drawback of the multi-year crop data training method is its high reliance on labelled data. To effectively train deep learning models, a large number of accurately labelled samples are needed. However, collecting such a vast amount of labelled data from diverse regions and years is a challenging task. The process is time-consuming, costly, and often impractical due to the effort and resources required. Moreover, it fails to address the issue of the cross-region issue.
To tackle the challenges of cross-year and cross-region adaptability, current methods adopt two main strategies:
From a sample perspective, fine-tuning pre-trained models with a few high-quality target domain samples is a common practice. This method adapts the original model to the new data distribution. For example, [26] improved deep models for nationwide land cover classification using high-confidence pseudo-labels. Similarly, [27] involved annotating new target domain samples to refine RF classifiers via active learning. However, this technique often necessitates labelling samples, which is not feasible for large-scale studies.
From a feature perspective, in the field of Unsupervised Domain Adaptation (UDA), earlier studies have developed DANN-based methods to align samples from various regions into a unified feature space, minimizing disparities in deep features. For instance, [17] introduced the STDAN, a novel unsupervised domain adaptation framework for crop type classification. Moreover, the DACCN [18] and the PAN [19], extended the loss function using the Maximum Mean Discrepancy (MMD) and the Multiple Kernel variant of Maximum Mean Discrepancy (MK-MMD), achieving improved accuracy compared to CNN and LSTM methods without domain adaptation. These methods often presuppose minor difference between the target and source domain data distributions. However, when these differences are substantial, the effectiveness diminishes.
This paper introduces the CropSTGAN framework, designed to tackle the challenge of cross-domain early crop mapping without the need for labelled data from the target domain, even when there are significant differences between the data distributions of the target and source domains. Unlike DANN-variant methods that focus on extracting invariant features, CropSTGAN transforms unlabeled target domain data into the source domain. The CropSTGAN system comprises three main components: a pre-processor, a CropSTGAN domain mapper, and a TempCNN crop mapper. The domain mapper, a key part of CropSTGAN, transforms time-series MSI data prepared by the pre-processor from the target domain into the source domain. Subsequently, a TempCNN crop mapper, pre-trained with labelled data from the source domain, processes this transformed data to produce the crop mapping results for the target domain. As far as our knowledge extends, we are the pioneering contributors to tackling the cross-domain issue in early crop mapping through the utilization of a model based on GAN combined with specially designed identity losses.
@@ III-A Study area
Figure 1: The Study Areas. (a) Jackson County, Minnesota, USA. (b) The Study Area (from 44.97N to 45.14N latitude and from 125.49W to 125.86W longitude), Jilin Province, China.
@@
Figure 2: The Average Time-series NDVI of Corn for the Source Domain and Target Domain. (a) Jackson County, 2019. (b) Jackson County, 2020. (c) Jackson County, 2021. (d) The study area of China, 2019.
This paper chose corn as the example crop to demonstrate the cross-domain capability of the proposed approach. The objective is to map corn cultivation locations in a target domain using models trained by labelled data from a source domain. The target and source domains could differ in their geographic locations (i.e., cross-region) or be in the same region in different years (i.e., cross-year). Note that, the proposed approach can be applied to other types of crops. Toward this end, the study areas include the two regions: Jackon County of the USA, and one study area from Jilin Province, China. For a visual depiction of the geographical locations of these areas, please refer to Figure 1.
TABLE I: The environmental conditions of the source domain and the target domains. ”SA” represents the Study Area. ”T” represents the yearly average temperature. ”P” stands for the average hourly precipitation. ”E” signifies the average hourly evaporation. ”R” indicates the surface net solar radiation.
SA Year T(K) P(mm/h) E(mm/h) R(kJ/m^2)
Jackson 2019 289.02 2.60 -1.01 4688.12
Jackson 2020 289.93 1.48 -1.11 5118.46
Jackson 2021 290.76 1.06 -1.06 5245.53
SA in China 2019 290.03 1.99 -1.99 12542.54
The study areas are characterized by unique environmental conditions such as temperature, precipitation, elevation, and solar radiation, which also vary over years. Table I presents the environmental metrics for these areas. Additionally, the corn cropping schedules differ among the regions, as depicted in Figure 3. These variations influence corn growth, resulting in distinct MSI feature patterns of the same crop across the different areas and time periods. Figures 2(a), 2(b), and 2(c) show the average NDVI values curves for the early growth stages of corn and other crops in Jackson County from 2019 to 2021. Figure 2(d) presents the curves in the study area of China. The corn growth NDVI patterns in Jackson County vary across the different years. Additionally, a clear distinction is evident between the NDVI values in Jackson County and those in the Chinese study area.